when x=3 and y=5, by how much does the value of 3x to the second power minus 2y exceed the value of 2x to the second power minus 3y...please explain i am having problems...thanks
Answer by Alan3354(69443)

3x to the second power minus 2y exceed the value of 2x to the second power minus 3y
--------
At x=3 and y=5:
3x^2-2y = 27-10 = 17
2x^2-3y = 18-15 = 3
17-3 = 14